# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

### v1

@@ VII-B Simulation Platform Integration
Fragmented scenario formats with unique syntax are adopted in the reviewed studies. Such heterogeneity often requires reworking prompts, templates, or conversion scripts. For example, studies [44, 53, 42] demonstrate how the same LLM (GPT-4) must be re-prompted to accommodate format differences.
@@
VII-C Lack of Domain-Specific Models
@@
General-purpose LLMs face notable limitations in scenario-based ADSs testing due to insufficient domain adaptation. This issue is exacerbated by the predominant reliance on prompt engineering over domain-specific fine-tuning, as well as the absence of authoritative ADSs datasets. For example, LLaMA generates invalid OpenSCENARIO syntax [53], while GPT-4 fails to interpret updated SUMO rules [44], revealing gaps in both semantic precision and temporal awareness.
@@
These limitations propagate across testing pipelines: biased scenario selection [51], syntax validation failure (e.g., ISO 21448 misinterpretation [36, 45]), and unclear performance evaluation in safety-critical contexts [41]. A hybrid approach combining domain fine-tuning and RAG could mitigate these issues by linking LLMs to structured knowledge.
VII-D Insufficient Attention to Scenario Database
@@
As discussed in previous sections, no studies have employed LLMs to construct scenario databases, which lags significantly behind other testing phases. This indicates a lack of systematic mechanisms for storing, managing, and reusing large volumes of generated scenarios. The construction of scenario databases should prioritize the development of standardized interfaces that integrate diverse data sources and convert them into machine-readable formats [19].
LLMs could assist in verifying the syntax and physical consistency of newly generated scenarios along with similarities, and performing de-duplication and quality assurance. Furthermore, LLM may support scenario analysis by comparing with regulations to generate coverage report.
VII-E Industrial Application Gap
@@
While LLMs show promise for scenario-based testing, their industrial adoption remains limited due to several constraints: 1) Data privacy and security: Utilizing LLMs often requires transmitting sensitive data (e.g., vehicle trajectories) to third-party cloud services (e.g., OpenAI Enterprise444https://openai.com/chatgpt/enterprise/), which risks violating local data protection regulations. Although enterprise APIs provide contractual safeguards, they typically prohibit model fine-tuning, thus restricting task-specific performance. Alternatively, deploying open-source LLMs locally can preserve data privacy, but requires substantial infrastructure investment. 2) Explainability: According to ISO 21448 [36], scenario generation must be traceable. However, LLMs often lack transparency in their generation process and cannot justify why specific scenarios are produced, thereby creating gaps in safety-critical applications.
VIII Conclusion
@@
This survey presents a structured overview of the application of LLMs in scenario-based testing of ADSs. By introducing a phase-based taxonomy, existing studies are systematically categorized according to roles LLM played in performed tasks in the testing pipeline. In addition, LLM types, usage strategies, and integrated simulators are quantitatively summarized, five key challenges that hinder LLM-enabled ADSs testing are identified.

### v2

@@ VI-B Simulation Platform Integration
Fragmented scenario formats with unique syntax are adopted in the reviewed studies. Such heterogeneity often requires reworking prompts, templates, or conversion scripts. For example, studies [Guzay.2023, Wang.2023, Lu.2024d] demonstrate how the same LLM (GPT-4) must be re-prompted to accommodate format differences.
@@
VI-C Lack of Domain-Specific Models
@@
General-purpose LLMs face notable limitations in scenario-based testing of ADSs due to insufficient domain adaptation. This issue is exacerbated by the predominant reliance on prompt engineering over domain-specific fine-tuning, as well as the absence of authoritative ADSs datasets. For example, LLaMA generates invalid OpenSCENARIO syntax [Wang.2023], while GPT-4 fails to interpret updated SUMO rules [Guzay.2023], revealing gaps in both semantic precision and temporal awareness.
@@
These limitations propagate across testing pipelines: biased scenario selection [Wu.2024], syntax validation failure (e.g., ISO 21448 misinterpretation [ISO21448, nouri2025simulation]), and unclear performance evaluation in safety-critical contexts [you2025comprehensive]. A hybrid approach combining domain fine-tuning and RAG could mitigate these issues by linking LLMs to structured knowledge.
Future work could focus on developing datasets that pair semantic descriptions with data sources or driving scenarios, enriched with diverse linguistic expressions for robustness. Domain-specific adaptation of LLMs through fine-tuning is needed to improve their handling of traffic rules, test scenario syntax, and temporal interactions. Finally, establishing dedicated benchmarks with metrics such as semantic alignment, and physical plausibility is crucial for systematic evaluation.
VI-D Insufficient Attention to Scenario Database
@@
As discussed in previous sections, no studies have employed LLMs to construct scenario databases, which lags significantly behind other testing phases. This indicates a lack of systematic mechanisms for storing, managing, and reusing large volumes of generated scenarios.
LLMs could contribute in three key directions. First, they can assist in verifying the syntax and physical consistency of newly generated scenarios along with similarities, and performing de-duplication and quality assurance. Second, they can facilitate structuring and similarity analysis, clustering functional scenarios and generating common descriptions of traffic and environment. For example, initiatives such as the “Language-to-Interaction” dataset [Yang.2023] demonstrates the feasibility of mapping expert knowledge into structured functional scenarios, thereby improving efficiency and coverage. Third, they can provide annotation and retrieval support by generating NL captions for logical and concrete scenarios, which serve both as semantic labels for selection and as training resources for domain-specific LLMs.
Although EU projects such as SUNRISE [sunrise-ccam], SYNERGIES [synergies-ccam], CERTAIN [certain-ccam] emphasize harmonization of scenario-based testing, LLM integration into scenario databases has not been considered. A key challenge is the lack of explainability, which constrains trust and adoption. Future research should therefore investigate systematic ways to embed LLMs into database construction while ensuring transparency and reliability.
VI-E Industrial Application Gap
@@
While LLMs show promise for scenario-based testing, their industrial adoption remains limited due to several constraints: 1) Data privacy and security: Utilizing LLMs often requires transmitting sensitive data (e.g., vehicle trajectories) to third-party cloud services (e.g., OpenAI Enterprise888https://openai.com/chatgpt/enterprise/), which risks violating local data protection regulations. Although enterprise APIs provide contractual safeguards, they typically prohibit model fine-tuning, thus restricting task-specific performance. Alternatively, deploying open-source LLMs locally can preserve data privacy, but requires substantial infrastructure investment. 2) Explainability and certification challenges: According to ISO 21448 [ISO21448], scenario generation must be traceable. However, LLMs often lack transparency in their generation process and cannot justify why specific scenarios are produced. This lack of explainability not only impedes industrial deployment but also creates obstacles for certification and regulatory approval, where authorities require rigorous traceability and justification of safety-critical decisions. Without addressing this gap, the integration of LLMs into certified testing pipelines remains highly challenging.
VII Conclusion
@@
This survey presents a structured overview of the application of LLMs in scenario-based testing of ADSs. By introducing a phase-based taxonomy, existing studies are systematically categorized according to roles that LLMs play across the testing pipeline. In addition, LLM types, usage strategies, and integrated simulators are quantitatively summarized, five key challenges that hinder LLM-enabled ADSs testing are identified.
@@
Yongqi Zhao received the bachelor’s degree from the China University of Petroleum (East China), Qingdao, China, in 2019, and the master’s degree from Technical University of Braunschweig, Braunschweig, Germany, in 2022. He is currently pursuing the Ph.D. degree with the Institute of Automotive Engineering, Graz University of Technology, Graz, Austria, with a research focus on virtual testing of automated driving systems. While pursuing his master’s degree, he gained practical experience through internships with Momenta, Stuttgart, Germany, and Volkswagen Group, Wolfsburg, Germany. |
Ji Zhou obtained the master’s degree (M.Sc) in Automotive Engineering from Jilin University, Changchun, China, in 2009. He is currently pursuing a Ph.D. degree at the Institute of Automotive Engineering, Graz University of Technology, Graz, Austria. He has worked in Ricardo, former PSA group and now in Stellantis group for more than 15 years. He has rich technical experience in powertrain control Software development, validation & calibration, vehicle Electrical & Electronic (EE) architecture integration validation, and EE product industrialization. His current research interests are mainly in the advanced methodology of Automated Driving Systems integration and validation. He has published 4 technical papers indexed by EI Compendex, 2 patents officially granted, and 11 other patents currently under publication. |
Dong Bi received the M.Sc. degree in Mechanical Engineering from Hubei University of Automotive Technology, Shiyan, China, in 2017. From 2017 to 2024, he served as a Lecturer at the same university, where he was involved in research and teaching in the areas of advanced driver assistance systems, next-generation electrical/electronic architecture (EEA), and functional safety testing for automated driving systems. Since October 2024, he has been pursuing the Ph.D. degree in the field of automated driving at Graz University of Technology, Graz, Austria. He has authored 6 academic papers and holds 5 invention patents. |
Tomislav Mihalj received his degree in Mechanical Engineering from the University of Zagreb, Croatia, in 2014, and earned his PhD from Graz University of Technology, Austria, in 2024. From 2014 to 2019, he worked as a Research Engineer at Virtual Vehicle, Graz, Austria, where he focused on the mechanical efficiency of combustion engines, vibration analysis, and crack propagation in wheel-rail contact. Between 2019 and 2024, he served as a University Project Assistant at the Institute of Automotive Engineering, Graz University of Technology, concentrating on the virtual verification of automated driving systems. Since 2024, he has been a Postdoctoral Researcher at the same institute, where he focuses on the verification and validation of driver assistance systems and supervises related research projects. He has authored or co-authored 12 peer-reviewed publications on rail vehicles and driver assistance systems. He has also contributed to several national and EU-funded projects related to automated driving. |
Jia Hu (Senior Member, IEEE) is currently the Zhongte Distinguished Chair of cooperative automation with the College of Transportation Engineering, Tongji University. Before joining Tongji University, he was a Research Associate with the Federal Highway Administration (FHWA), USA. He is a member of the TRB (Division of the National Academies) Vehicle Highway Automation Committee, the Freeway Operation Committee, and the Simulation Subcommittee of Traffic Signal Systems Committee, and a member of the CAV Impact Committee and the Artificial Intelligence Committee of the ASCE Transportation and Development Institute. He is an Advisory Editorial Board Member of Transportation Research Part C: Emerging Technologies. He has been an Associate Editor of the IEEE Intelligent Vehicles Symposium since 2018 and the IEEE Intelligent Transportation Systems Conference since 2019. He is an Associate Editor of Journal of Transportation Engineering (American Society of Civil Engineers) and Journal of Intelligent Transportation Systems. |
Arno Eichberger (Member, IEEE) received the degree in mechanical engineering and the Ph.D. degree (Hons.) in technical sciences from the Graz University of Technology, Graz, Austria, in 1995 and 1998, respectively. From 1998 to 2007, he was employed with Magna Steyr Fahrzeugtechnik AG&Company, Graz, where he dealt with different aspects of active and passive safety. Since 2007, he has been working with the Institute of Automotive Engineering, Graz University of Technology, dealing with driver assistance systems, vehicle dynamics, and suspensions. Since 2012, he has been an Associate Professor holding a “venia docendi” of automotive engineering. |
![[Uncaptioned image]](biofig/Yongqi_Zhao.png)
![[Uncaptioned image]](biofig/ji_zhou.png)
![[Uncaptioned image]](biofig/dong_bi.png)
![[Uncaptioned image]](biofig/tomislav_mihalj.png)
![[Uncaptioned image]](biofig/jia_hu.png)
![[Uncaptioned image]](biofig/arno_eichberger.png)